A 7kg object is at the top of a 3m tall ramp.Due to friction it is only moving 5m/s at the bottom.Calculate efficiency

g=9.8
1
Expert's answer
2020-10-29T07:00:26-0400
"\eta=\frac{0.5mv^2}{mgh}\\\eta=\frac{0.5(5)^2}{(9.8)(3)}=0.43"